OpenStudy (sedatefrog712):

Could someone show me how to solve this step by step (cos Θ + cos Θ)^2 + (cos Θ + cos Θ)^2

OpenStudy (sedatefrog712):

@PRAETORIAN.10

OpenStudy (sedatefrog712):

@UnkleRhaukus

OpenStudy (unklerhaukus):

\[(\cos\theta + \cos\theta)^2 + (\cos\theta + \cos\theta)^2\\ =(2\cos\theta)^2 + (2\cos\theta)^2\\ = 2(2\cos\theta)^2\\ = 2(4\cos^2\theta)\\ = 8\cos^2\theta\]

OpenStudy (sedatefrog712):

thank you so much you are a legend

OpenStudy (sedatefrog712):

now did you use foil for this

OpenStudy (unklerhaukus):

i did not use foil,

OpenStudy (sedatefrog712):

ok one last question how did you get 2(4cos^2)theta

OpenStudy (unklerhaukus):

\[(\cos\theta + \cos\theta)^2 + (\cos\theta + \cos\theta)^2\\ =(2\cos\theta)^2 + (2\cos\theta)^2\\ = 2(2\cos\theta)^2\\ = 2(2^2\cos^2\theta)\\ = 2(4\cos^2\theta)\\ = 8\cos^2\theta\]

OpenStudy (unklerhaukus):

Using foil:\[(\cos\theta + \cos\theta)^2 + (\cos\theta + \cos\theta)^2\\ =(\cos^2\theta+\cos^2\theta +\cos^2\theta+\cos^2\theta) + (\cos^2\theta+\cos^2\theta+\cos^2\theta+\cos^2\theta)\\ = (4\cos^2\theta)+(4\cos^2\theta)\\ = 8\cos^2\theta\]

OpenStudy (sedatefrog712):

oh ok that makes sense thank you so much im going to add this to my notebook :D